OpenStudy (anonymous):

The measure of an angle is 44 more than the measure of its supplement. Find the measure of the angles.

OpenStudy (anonymous):

let the angles be x and x+44 x+(x+44)=180 x+x+44=180 2x=180-44 2x=136 x=136/2 x=68 therefore, angles are 68 and 68+44=112